# Principled Design of Interpretable Automated Scoring for Large-Scale Educational Assessments

### v1

@@ Abstract
AI-driven automated scoring systems offer scalable and efficient means of evaluating complex student-generated responses. Yet, despite increasing demand for transparency and interpretability, the field has yet to develop a widely accepted solution for interpretable automated scoring to be used in large-scale real-world assessments. This work takes a principled approach to address this challenge. We analyze the needs and potential benefits of interpretable automated scoring for various assessment stakeholders and develop four principles of interpretability – Faithfulness, Groundedness, Traceability, and Interchangeability (FGTI) – targeted at those needs. To illustrate the feasibility of implementing these principles, we develop the AnalyticScore framework for short answer scoring as a baseline reference framework for future research. AnalyticScore operates by (1) extracting explicitly identifiable elements of the responses, (2) featurizing each response into human-interpretable values using LLMs, and (3) applying an intuitive ordinal logistic regression model for scoring. In terms of scoring accuracy, AnalyticScore outperforms many uninterpretable scoring methods, and is within only 0.06 QWK of the uninterpretable SOTA on average across 10 items from the ASAP-SAS dataset. By comparing against human annotators conducting the same featurization task, we further demonstrate that the featurization behavior of AnalyticScore aligns well with that of humans. footnotetext: * Equal Advising
@@ 1 Introduction
Accurate and credible assessment of knowledge and skills forms the basis for effective decision making in a variety of educational contexts, from student learning and instructional design to program development and policy making Berman et al. (2019). When the set of knowledge and skills to be gauged involves complex, open-ended problem-solving and communication abilities, AI-driven automated scoring systems can offer rapid, accessible, and scalable alternatives to the otherwise labor-intensive and costly process of training and deploying human scorers Foltz et al. (2020). Automated scoring systems have been increasingly adopted across various assessment contexts over the past several decades. Today’s scoring algorithms achieve acceptable levels of scoring accuracy in various areas of human learning Whitmer and Beiting-Parrish (2023, 2024).
@@
In this paper, we take a principled approach towards building a practical interpretable automated scoring solution for large-scale assessments. An effective interpretability solution begins by identifying the diverse needs of each stakeholder in understanding the system’s decisions, and by grounding the development of interpretable AI systems in those needs Bhatt et al. (2020), Preece et al. (2018), Páez (2019). Research on explainable automated scoring, on the other hand, has largely ignored this need-finding process. As we observe later in this paper (Section 2), this neglect has often led to several claimed interpretability solutions that fail to address the diverse and nuanced interpretability needs of the human actors in educational assessment.
We identify the needs and benefits of model explanations for various large-scale assessment stakeholder groups consisting of test takers, assessment developers, and test users (Section 2.1). Targeted at those needs, we develop the principles of faithful, grounded, traceable, and interchangeable model interpretations for AI-driven automated scoring (Section 2.2).
We further illustrate the feasibility of implementing these principles in practice and establish a concrete baseline for future work (Section 3). AnalyticScore is the first interpretable automated short-answer scoring framework to embody our principles. It operates by extracting explicitly identifiable elements from unannotated response texts and featurizing each response into human-interpretable values based on those elements. These features are input to an intuitive ordinal logistic regression module for scoring.
We measure the performance of AnalyticScore on a real-world response dataset by measuring (1) scoring accuracy and (2) alignment of featurization behaviors with human judgments (Sections 4 and 5). AnalyticScore outperforms many uninterpretable scoring methods, and is within only 0.06 QWK of the uninterpretable SOTA on average across 10 items from the ASAP-SAS dataset. The featurization behavior of AnalyticScore also aligns well with humans (0.90, 0.72, 0.81 QWK across assessment areas). Our findings indicate strong potential for implementing accurate and well-aligned interpretability solutions that meet the real needs of assessment stakeholders.
@@
As AI-driven automated scoring systems became more complex and opaque, researchers have increasingly noted the need to enhance the transparency of these systems through model explanations Bennett and Zhang (2015), Bauer and Zapata-Rivera (2020), Schlippe et al. (2022). Several approaches have been proposed to address interpretable automated scoring, and we discuss them in detail in Section 2.2 in connection with our four principles. Despite growing research interests, interpretable automated scoring still lacks practical adoption and meaningful field use. The 2023 NAEP Math Automated Scoring Challenge [^1] https://github.com/NAEP-AS-Challenge/math-prediction for open-ended math responses organized by the US National Center for Education Statistics (NCES) found that none of the submissions met the criteria for “interpretability” despite several methods achieving near-human scoring accuracy Institute of Education Statistics (2023), Whitmer and Beiting-Parrish (2024). This gap highlights the need for human-centered explainability solutions driven by assessment stakeholders’ real needs.
@@ 2 Building the Principles of Interpretable Automated Scoring
Insights derived from scoring support various stakeholders throughout the overall assessment process. Below we analyze three main stakeholder groups in large-scale assessment – test takers, assessment developers, and test users Berman et al. (2019), AERA et al. (2014). Each stakeholder group’s distinct roles and priorities uniquely shape how interpretable automated scoring can address their specific needs and improve their assessment experience.
2.1 Explainability Needs and Potential Benefits
@@
Test users, including professionals who select and administer tests, educators, administrators, and policymakers, depend on score reliability and interpretation validity to make system-level decisions or instructional differentiation. Their reliance on the integrity and validity of scores to drive decisions is significant AERA et al. (2014). Model explanations provide concrete evidence to validate the choice of the scoring model [^2] More examples of validity arguments on the use of automated scoring can be found in (Bennett and Zhang, 2015, Table 7.7).. This includes understanding whether the extracted features and scoring logic fully capture the rubric and the construct definition, and whether the internal structure of the automated scores align with the construct of interest Bennett and Zhang (2015).
@@ 2.2 The FGTI Interpretability Principles
A notable example of unfaithful scoring explanations are texts produced by prompting LLMs to generate an explanation (e.g., Lee et al. (2024) and Li et al. (2025)). Stepwise reasoning verbalized by LLM through prompting strategies such as chain-of-thought Wei et al. (2022) are not explanations of their internal computation Sarkar (2024) and often fail to reflect the model’s true reasoning behavior Turpin et al. (2023), Arcuschin et al. (2025). Moreover, LLMs are highly sensitive to superficial changes in prompts and input text, frequently exhibiting inconsistent judgments Wang et al. (2024). Therefore, even when LLMs achieve high scoring accuracy, prompting them for explanations cannot reliably address the stakeholder needs identified in Section 2.1.
@@
How should the model process these features to ultimately produce a final score? Scoring is inherently an evidentiary reasoning process, where elements of the student response and item tasks serve as evidence to support the inference about student knowledge and skills that the score represents DiCerbo et al. (2020), Mislevy (2020). Stakeholders need to be able to inspect and interact with the internal structure of the scoring model to ensure soundness, construct-relevance, fairness, and model understanding (Section 2.1). To meet this need, the model’s evidentiary reasoning process must be decomposable into clear, sequential steps that a human could reliably execute and possibly intervene. Our next 2 principles state that the scoring model should be conducive to this decomposition and intervention:
@@
Not all intermediate representations calculated by the model need to be understandable by human, but the reasoning subroutines should collectively account for the entire scoring logic. Moreover, humans should be able to act interchangeably with each decomposed module and replace the module outputs with human-generated results if deemed necessary.
Many proposed interpretability approaches are not grounded, traceable, or interchangeable in the sense described above. These include, for instance, calculating feature importance values Kumar and Boulanger (2021, 2020), Schlippe et al. (2022), Asazuma et al. (2023), displaying feature attribution maps Schlippe et al. (2022), Li et al. (2025), or presenting confidence metrics for scoring decisions Conijn et al. (2023). This limits the capacity to thoroughly inspect the model’s features and internal structure, which is critical to meeting the needs of the stakeholders.
@@ 3 A Principled Framework for Interpretable Automated Scoring
How would the FGTI principles be implemented in practice? To illustrate the feasibility of implementing these principles and to set a baseline for future research, we present AnalyticScore as a reference framework in the domain of short-answer scoring. In this setting, students write a short 1-5 sentence answer in response to an assessment item which is scored with an emphasis on content correctness and demonstrated reasoning Leacock and Chodorow (2003), Shermis (2015). The scoring model has access to a training set of student response texts paired with human-annotated scores $(r_{1},s_{1}),...,(r_{n},s_{n})$ and possibly additional unannotated responses $\{r_{n+1},...,r_{m}\}$. The goal at inference time is to predict the score $s$ for a new response $r$.
AnalyticScore (Figure 1) is a 3-phase, LLM-based framework grounded in our four principles of interpretable automated scoring. Phase 1 identifies explicitly grounded analytic components to be used. Phase 2 catalogs, or featurizes, the presence of these components in student responses. Phase 3 uses the features to compute a score. Phases 1 and 2 depend only on the response texts without any annotations. Human score labels are only used with Phase 3.
3.1 Phase 1: Extracting Analytic Components.
With the response texts from the training set (and optionally assessment content), AnalyticScore first extracts a set of analytic components, which are explicitly identifiable elements of student responses described in Principle 2. In this work, we consider a specific type of components which are representative, atomic units of explicit statements, arguments, or claims as in Figure 1.
@@
Component extraction is implemented using an LLM with the prompts shown in Figure 2. Having too many analytic components could diminish the interchangeability (Principle 4) of the overall framework by exploding the number of features used in scoring Lipton (2018), so we limit to generating 15 components per request.
@@ 3.2 Phase 2: Featurizing Responses
Once the analytic components have been identified, student responses are featurized according to the presence of these components $c_{1},...,c_{m}$ in each response $r$. This step uses a labeling function $f(r;c)$ whose outputs are associated with human-understandable meaning (Principles  2 &  3). The exact label definitions used can be selected using natural language. In this work, we explore the following general purpose labeling function for $f(r;c)$:
@@
We implement $f(r;c)$ using a Chain-of-Thought Wei et al. (2022) prompting template shown in Figure 2333Note that we CoT is used solely as a prompting technique, and the generated “thoughts” are explicitly discarded. . Inspired by the self-consistency decoding strategy for LLMs Wang et al. (2022), we apply the first-to-three aggregation rule to consider the possibly diverse interpretation of the labeling criteria when selecting the final output. Easily interpretable one-hot encodings of each $f(r;c_{m})$ are then concatenated to produce a $3m$-dimensional binary featurization of $r$:
$\text{Featurize}(r)=\text{OneHot}(f(r;c_{1}))\operatorname*{\mathbin{\|}}\cdots\operatorname*{\mathbin{\|}}\text{OneHot}(f(r;c_{k}))$
Figure 2: Prompts used in AnalyticScore
@@
Using proprietary LLMs for featurization can quickly become too expensive in large-scale assessment settings, especially with many analytic components. To avoid the linearly growing cost of featurization, we supervised fine-tuned a small open-source model using a subsample of $(r,c)$ pairs, where $r$ is a response from the training set and $c$ is an analytic component from Phase 1. More specifically, we randomly sampled 10k pairs across all 10 items, calculated the featurization labels on these samples using o4-mini, and collected the full LLM model requests and outputs generated during this process which aligned with the aggregated final decision. This dataset was used to fine-tune Llama-3.1-8b-instruct with QLoRA Dettmers et al. (2023).
@@ 3.4 Analysis of AnalyticScore
An example of AnalyticScore’s model explanation is shown in the right panel of Figure 1. By demonstrating human-understandable features of the response (Principle 2) and the exact decision process (Principle 3), the explanation transparently and faithfully reveals the actual scoring mechanism used (Principle 1). If, based on the explanation, the model is suspected to have made an error (e.g., C6 should be a check, not a triangle), a human inspector can modify the featurization and rerun the scoring algorithm (Principle 4), which is also how the “if instead…” explanation is generated.
The structure of AnalyticScore’s scoring model is akin to Concept Bottleneck Models Koh et al. (2020), Yang et al. (2023) in that we enforce a layer of intermediate representations with human-understandable “concepts.” Our framework ensures that the intermediate features have human-understandable values that are associated with explicitly identifiable elements (Principle 2), as opposed to characteristics that are inferred from the response.
@@ 4 Evaluating AnalyticScore
Having introduced AnalyticScore and discussed its interpretability, we now evaluate its scoring performance and how its featurization aligns with human judgments on a real-world response scoring dataset.
Token Len. Train Valid Test Assessment Area
Q1 $47.5\pm 22.2$ 1,341 331 557 Science
Q2 $59.2\pm 22.6$ 1,024 254 426
Q3 $47.9\pm 14.6$ 1,445 363 406
Reading
(Informational Text)
Q4 $40.3\pm 15.5$ 1,308 349 295
Q5 $25.1\pm 21.5$ 1,459 336 598 Science
Q6 $23.8\pm 22.6$ 1,418 379 599
Q7 $41.3\pm 25.1$ 1,432 367 599
Reading
Q8 $53.0\pm 32.6$ 1,446 353 599
Q9 $49.7\pm 36.3$ 1,453 345 599
Reading
(Informational Text)
Q10 $41.1\pm 28.5$ 1,314 326 546 Science
Table 1: ASAP-SAS dataset detail by item.
@@
The ASAP-SAS dataset Shermis (2015) [^4] https://www.kaggle.com/competitions/asap-sas/data is the largest publicly available dataset with short answer responses from schoolchildren for 10 different open-ended exam questions. Human raters double-scored and assigned a single number to each student response using a 3 or 4 point rubric. The assessment area for each question, as well as the sample sizes and response lengths are reported in Table 1. We use the original test set and split the public training set into training and validation sets with a 8:2 ratio.
@@
For each assessment item, we used GPT-4.1 as the base LLM and extracted 15 analytic components except for Q7. This item uses a two-part scoring scheme to separately assess a character trait identified from the reading and its supporting evidence. We extracted 15 analytic components from each part, totaling 30 components. For the featurizer, we experimented with GPT-4.1-mini and Llama-3.1-8B-Instruct as our base LLM, each with temperature set to 0.7 and 1.0. We distilled the Llama featurizer for 2 epochs using a batch size of 4 and learning rate of 1e-4. All model calls were made through the official OpenAI API. Fine-tuning was conducted on an Ubuntu 20.04 machine with 2 RTX A6000 GPUs (49Gb memory), 16 AMD EPYC 9224 24-Core Processors, and 250Gb of CPU RAM.

### v2

@@ Abstract
AI-driven automated scoring systems offer scalable and efficient means of evaluating complex student-generated responses. Yet, despite increasing demand for transparency and interpretability, the field has yet to develop a widely accepted solution for interpretable automated scoring to be used in large-scale real-world assessments. This work takes a principled approach to address this challenge. We analyze the needs and potential benefits of interpretable automated scoring for various assessment stakeholder groups and develop four principles of interpretability – Faithfulness, Groundedness, Traceability, and Interchangeability (FGTI) – targeted at those needs. To illustrate the feasibility of implementing these principles, we develop the AnalyticScore framework for short answer scoring as a baseline reference framework for future research. In terms of scoring accuracy, AnalyticScore outperforms many uninterpretable scoring methods and is, on average, within 0.06 QWK of the uninterpretable SOTA across 10 items from the ASAP-SAS dataset. By comparing against human annotators conducting the same featurization task, we further demonstrate that the featurization behavior of AnalyticScore aligns well with that of humans.
footnotetext: * Equal Advising
@@ 1 Introduction
Accurate and credible assessment of knowledge and skills forms the basis for effective decision making in a variety of educational contexts, from student learning and instructional design to program development and policy making Berman et al. (2019). When the set of knowledge and skills to be gauged involves complex, open-ended problem-solving and communication abilities, AI-driven automated scoring systems can offer rapid, accessible, and scalable alternatives to the otherwise labor-intensive and costly process of training and deploying human scorers Foltz et al. (2020). Automated scoring systems have been increasingly adopted across many assessment contexts over the past several decades, achieving acceptable levels of scoring accuracy in various areas of human learning Whitmer and Beiting-Parrish (2023, 2024).
@@
In this paper, we take a principled approach towards building a practical interpretable automated scoring solution for large-scale assessments. An effective interpretability solution begins by identifying the diverse needs of each stakeholder in understanding the system’s decisions, and by grounding the development of interpretable AI systems in those needs Bhatt et al. (2020), Preece et al. (2018), Páez (2019). Research on interpretable automated scoring, on the other hand, has largely ignored this need-finding process. As we discuss in Section 2.2, this neglect has often led to several claimed interpretability solutions that fail to address the diverse and nuanced interpretability needs of the human actors in educational assessment.
We identify the needs and benefits of scoring model explanations for various large-scale assessment stakeholder groups consisting of test takers, assessment developers, and test users (Section 2.1). Targeted at those needs, we develop the principles of faithful, grounded, traceable, and interchangeable (FGTI) model interpretations for AI-driven automated scoring (Section 2.2).
We further illustrate the feasibility of implementing the FGTI principles in practice and establish a concrete baseline for future work (Section 3). AnalyticScore is the first interpretable automated short-answer scoring framework to embody our principles. It operates by extracting explicitly identifiable elements from unannotated response texts and featurizing each response into human-interpretable values based on those elements. These features are input to an intuitive ordinal logistic regression module for scoring.
We measure the performance of AnalyticScore on a real-world response dataset by measuring (1) scoring accuracy and (2) alignment of featurization behaviors with human judgments (Sections 4 and 5). AnalyticScore outperforms many uninterpretable scoring methods and is within 0.06 QWK of the uninterpretable SOTA on average across 10 items from the ASAP-SAS dataset. The featurization behavior of AnalyticScore also aligns well with humans (0.90, 0.72, 0.81 QWK across assessment areas). Our findings indicate strong potential for implementing accurate and well-aligned interpretability solutions that meet the real needs of assessment stakeholders.
@@
Researchers have increasingly noted the need to enhance the transparency of complex AI-driven automated scoring systems through model explanations Bennett and Zhang (2015), Bauer and Zapata-Rivera (2020), Schlippe et al. (2022). Several approaches have been proposed to address this challenge, and we discuss them in detail in Section 2.2 in connection with our four principles.
Despite growing research interests, interpretable automated scoring still lacks practical adoption and meaningful field use. The 2023 NAEP Math Automated Scoring Challenge [^1] https://github.com/NAEP-AS-Challenge/math-prediction for open-ended math responses organized by the US National Center for Education Statistics (NCES) found that none of the submissions met the criteria for interpretability despite several methods achieving near-human scoring accuracy Institute of Education Statistics (2023), Whitmer and Beiting-Parrish (2024). Our work provides a stakeholder-centered approach towards addressing this gap.
@@ 2 Building the Principles of Interpretable Automated Scoring
Insights derived from scoring support various human actors throughout the overall assessment process. Below we analyze three main stakeholder groups in large-scale assessment – test takers, assessment developers, and test users Berman et al. (2019), AERA et al. (2014). Each stakeholder group’s distinct roles and priorities uniquely shape how interpretable automated scoring can improve their assessment experience.
2.1 Interpretability Needs and Benefits
@@
Test users, including professionals who select and administer tests, educators, administrators, and policymakers, depend on score reliability and interpretation validity to make system-level or instructional decisions. Their reliance on the integrity and validity of scores to drive decisions is significant AERA et al. (2014). Model explanations provide concrete evidence to validate the choice of the scoring model [^2] More examples of validity arguments on the use of automated scoring can be found in (Bennett and Zhang, 2015, Table 7.7).. This includes understanding whether the extracted features and scoring logic fully capture the rubric and the construct definition, and whether the internal structure of the automated scores align with the construct of interest Bennett and Zhang (2015).
@@ 2.2 The FGTI Interpretability Principles
A notable example of unfaithful scoring explanations are texts produced by prompting LLMs to generate an explanation (e.g., Lee et al. (2024) and Li et al. (2025)). Stepwise reasoning verbalized by LLM through prompting strategies such as chain-of-thought Wei et al. (2022) are not explanations of their internal computation Sarkar (2024) and often fail to reflect the model’s true reasoning behavior Turpin et al. (2023), Arcuschin et al. (2025). Moreover, LLMs are highly sensitive to superficial changes in prompts and input text, frequently exhibiting inconsistent judgments Wang et al. (2024). Therefore, even when LLMs achieve high scoring accuracy, prompting them to “explain their decisions” cannot reliably address the stakeholder needs identified in Section 2.1.
@@
How should the model process these features to ultimately produce a final score? Scoring is inherently an evidentiary reasoning process, where elements of the student response and item tasks serve as evidence to support the inference about student knowledge and skills that the score represents DiCerbo et al. (2020), Mislevy (2020). Stakeholders need to be able to inspect and interact with the internal structure of the scoring model to ensure soundness, construct-relevance, and fairness (Section 2.1). To meet this need, the model’s evidentiary reasoning process must be decomposable into clear, sequential steps that a human could reliably execute and possibly intervene. Our next 2 principles state that the scoring model should be conducive to this decomposition and intervention:
@@
Not all intermediate representations calculated by the model need to be understandable by human, but the reasoning subroutines should collectively account for the entire scoring logic. Moreover, humans should be able to act interchangeably with each subroutine and replace its outputs with human-generated results if deemed necessary.
Many proposed interpretability approaches are not grounded, traceable, or interchangeable in the sense described above. These include, for instance, calculating feature importance values Kumar and Boulanger (2021, 2020), Schlippe et al. (2022), Asazuma et al. (2023), displaying feature attribution maps Schlippe et al. (2022), Li et al. (2025), or presenting confidence metrics for scoring decisions Conijn et al. (2023). This limits the capacity to thoroughly inspect the model’s features and internal decision-making structure.
@@ 3 A Principled Framework for Interpretable Automated Scoring
To illustrate the feasibility of implementing the FGTI principles and to set a baseline for future research, we present AnalyticScore as a reference framework in the domain of short-answer scoring. In this setting, students write a short 1-5 sentence answer in response to an assessment item which is scored with an emphasis on content correctness and demonstrated reasoning Leacock and Chodorow (2003), Shermis (2015). The scoring model has access to a training set of student response texts paired with human-annotated scores $(r_{1},s_{1}),...,(r_{n},s_{n})$ as well as unannotated responses $\{r_{n+1},...,r_{m}\}$. The goal at inference time is to predict the score $s$ for a new response $r$.
AnalyticScore (Figure 1) is a 3-phase, LLM-based framework embodying the FGTI principles. Phase 1 identifies explicitly grounded analytic components to be used. Phase 2 catalogs, or featurizes, the presence of these components in student responses. Phase 3 uses the features to compute a score. Phases 1 and 2 depend only on the response texts without any annotations. Human score labels are only used during Phase 3. Source code implementing AnalyticScore can be found at this link: https://github.com/yunsungkim0908/analyticscore.
3.1 Phase 1: Extracting Analytic Components
With the response texts from the training set (and optionally the content of the assessment item), AnalyticScore first extracts a set of analytic components, which are explicitly identifiable elements of student responses described in Principle 2. For the short-answer scoring task considered in this work, we operationalize analytic components as propositions.
@@ 3.1 Phase 1: Extracting Analytic Components
Provided the extracted components align with the intended item design and are human understandable, any procedure can be used for component extraction. If necessary, these components could also be manually reviewed and modified by assessment designers. In this work, we implement component extraction by prompting an LLM with the prompts shown in Appendix A. Having too many analytic components could reduce interchangeability (Principle 4) by increasing the number of processed features Lipton (2018). We therefore limit extraction to 15 components per scoring unit.
@@ 3.2 Phase 2: Featurizing Responses
Once the analytic components have been identified, student responses are featurized according to the presence of these components $c_{1},...,c_{k}$ in each response $r$. This step uses a labeling function $f(r;c)$ whose outputs are associated with human-understandable meaning (Principles  2 &  3). The exact label definitions used can be selected using natural language. In this work, we explore the following general purpose labeling function for $f(r;c)$. The precise definition of each score category can be found in Appendix A:
@@
We implement $f(r;c)$ using Chain-of-Thought Wei et al. (2022) prompting using the prompts shown in Appendix A. (Note that CoT is used solely as a prompting technique, and the generated “thoughts” are explicitly discarded.) Inspired by the self-consistency decoding strategy for LLMs Wang et al. (2022), we apply the first-to-three aggregation rule to consider the possibly diverse interpretation of the labeling criteria when selecting the final output. Easily interpretable one-hot encodings of each $f(r;c)$ are then concatenated to produce a $3k$-dimensional binary featurization of $r$:
$F(r)=\text{OneHot}(f(r;c_{1}))\operatorname*{\mathbin{\|}}\cdots\operatorname*{\mathbin{\|}}\text{OneHot}(f(r;c_{k}))$
@@
Using proprietary LLMs for featurization can quickly become too expensive in large-scale assessment settings, especially with many analytic components. To avoid the linearly growing cost of featurization, we supervised fine-tuned a small open-source model using a subsample of $(r,c)$ pairs, where $r$ is a response from the training set and $c$ is an analytic component from Phase 1. More specifically, we randomly sampled 10k pairs across all 10 items, calculated the featurization labels on these samples using o4-mini, and collected the LLM model requests and outputs generated during this process that aligned with the aggregated final decision. This dataset was used to fine-tune Llama-3.1-8b-instruct with QLoRA Dettmers et al. (2023).
@@ 3.4 Analysis of AnalyticScore
An example of AnalyticScore’s model explanation is shown in the far right panel of Figure 1. By demonstrating human-understandable features of the response (Principle 2) and the exact decision process (Principle 3), the explanation transparently and faithfully reveals the actual scoring mechanism used (Principle 1). If, based on the explanation, the model is suspected to have made an error (e.g., C6 should be a check, not a triangle), a human inspector can modify the featurization and recalculate the score by following the scoring algorithm (Principle 4), which is also how the “if instead…” explanation is generated.
The structure of AnalyticScore’s scoring model is akin to Concept Bottleneck Models Koh et al. (2020), Yang et al. (2023) in that we enforce a layer of intermediate representations with human-understandable “concepts.” Our framework ensures that the intermediate features have human-understandable values associated with explicitly identifiable elements of student work (Principle 2), as opposed to inferred characteristics.
@@ 4 Evaluating AnalyticScore
We now evaluate AnalyticScore’s scoring accuracy and how well its featurization aligns with human judgments on a real-world open-ended response dataset.
@@
The ASAP-SAS dataset Shermis (2015) [^3] https://www.kaggle.com/competitions/asap-sas/data is the largest publicly available dataset of English short answer responses from US schoolchildren for 10 different open-ended exam questions, covering 3 areas of assessment: Science, Reading (Informational Text), and Reading (Literature Text). Human raters double-scored and assigned a single number to each student response using a 3 or 4 point rubric. The dataset details are reported in Appendix B. We use the original test set and split the public training set into training and validation sets with an 8:2 ratio.
@@
For each assessment item, we used GPT-4.1 as the base LLM and extracted 15 analytic components except for Q7. This item uses a two-part scoring scheme to separately assess a character trait identified from the reading and its supporting evidence. We extracted 15 analytic components from each part, totaling 30 components. For the featurizer, we experimented with GPT-4.1-mini and Llama-3.1-8B-Instruct as our base LLM, each with temperature set to 0.7 and 1.0. Training and implementation details can be found in Appendix C.